OpenStudy (anonymous):

a.) If it is exponential or decay b.) the rate of growth or decay c.) the y-intercept 1.) f(x)=123(0.95)^x 2.) g(x)=75(1.07)^x 3.)h(x)=6(3)^x

OpenStudy (campbell_st):

(c) the y- intercept occurs when x = 0 because each term is a power of x and when x = 0 you have a^0 = 1 or more specifically 10(a)^0 = 10 * 1 = 10 hope this helps...